OpenStudy (anonymous):

Evaluate the expression. log_4 0.125

OpenStudy (anonymous):

Similar to the previous one, but a little trickier.\[\log_{4} (0.125)=\log_{4}(1/8)=\log_{4}(4^{-3/2})=-3/2\]

OpenStudy (anonymous):

This one depends on you knowing that 8 = 4*2 = 4*4^(1/2) = 4^(3/2).